# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

### v2

@@ B.2 Proof of Lemma 2
B.3 Proof of Lemma 3
We first upper bound the higher-order cumulants. Let $t=\frac{\rho}{3\sqrt{1+\rho^{2}}}$. Then $0<t<\frac{1}{1+\rho}$ since $10<\rho<$ 1. By Lemma 1,
@@ B.3 Proof of Lemma 3
It remains to upper bound $-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$. We note that
@@ B.4 Proof of Lemma 4
It remains to upper bound $-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$. We note that
@@
Appendix C Auxiliary results
@@
where the last inequality is due to the fact that $(1-x)\log(1-x)+x\geq 0$ for any $10<x<$ 1 and $0<\eta\leq\frac{\gamma}{4(1+\gamma)}<\frac{1}{4}$. Since $\eta\leq\frac{\gamma}{4(1+\gamma)}$ and $(1+\gamma)\log(1+\gamma)-\gamma\geq\frac{\gamma^{2}}{2(1+\gamma)}$, we obtain $\eta\gamma\leq\frac{\gamma^{2}}{4(1+\gamma)}\leq\frac{1}{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]$. Therefore,
@@
where $\mathrm{(a)}$ is because $h(x)=h(1-x)$ and $h(x)\geq-x\log x$; $\mathrm{(b)}$ is because $\log\left(\frac{m}{k}\right)\geq\log 2$ when $k\leq\frac{m}{2}$; $\mathrm{(c)}$ is because $k\log\left(\frac{m}{k}\right)\geq\log m$ for $1\leq k\leq 2\log m$ when $m\geq 10$; $\mathrm{(d)}$ is because $2\cdot 2^{-2\log m}\leq\frac{2}{m}$. ∎
@@
=$\overset{\mathrm{(b)}}{=}$ $~{}p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right],$
where $\mathrm{(a)}$ uses $\log(1+x)\leq x$ for any $x>-1$; $\mathrm{(b)}$ follows from $\gamma=\frac{\rho(1-p)}{p}$. Since $\log(1+x)\geq\frac{x}{x+1}+\frac{x^{2}}{2(x+1)^{2}}$ for any $x\geq 0$, we obtain that $p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]\geq\frac{p^{2}\gamma^{2}}{2(\gamma+1)}$. When $\gamma<3$, since $\frac{p^{2}\gamma^{2}}{2(\gamma+1)}\geq\frac{p^{2}\gamma^{2}}{8}=\frac{\rho^{2}\cdot 3(1-p)^{2}}{24}\geq\frac{\rho^{2}\left[2p(1-p)+p^{2}\right]}{24}$ for $0<p\leq\frac{1}{2}$, we obtain that

### v3

@@ C.2 Proof of Lemma 2
C.3 Proof of Lemma 3
We first upper bound the higher-order cumulants. Let $t=\frac{\rho}{3\sqrt{1+\rho^{2}}}$. Then $0<t<\frac{1}{1+\rho}$ since $0<\rho<1$. By Lemma 1,
@@ C.3 Proof of Lemma 3
It remains to upper bound $t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$. We note that
@@ C.4 Proof of Lemma 4
It remains to upper bound $t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$. We note that
@@
C.5 Proof of Lemma 5
The size of ${\mathcal{M}}_{\delta}$ follows from the standard volume argument [PW25, Theorem 27.3]. For $r\in[m]$, let $B(\pi,r)\triangleq\{\pi^{\prime}:d(\pi,\pi^{\prime})\leq r\}$ denote the ball of radius $r$ centered at $\pi$. Then, we obtain
It remains to evaluate the cardinality of $\mathcal{S}_{n,m}$ and upper bound the volume of the ball under our distance metric $d$. It is straightforward to obtain that $|\mathcal{S}_{n,m}|=\binom{n}{m}^{2}m!.$ Let $k=\delta m$. Note that all elements from $B(\pi,m-k)$ have at least $k$ common mappings. To upper bound $|B(\pi,m-k)|$, we first choose $k$ elements from the domain of $\pi$ and map to the same value as $\pi$, and the remaining domain and range of size mkm-k and the mapping are selected arbitrarily. We get $|B(\pi,m-k)|\leq\binom{m}{k}\binom{n-k}{m-k}^{2}(m-k)!.$ Consequently,
where we use the inequalities that $(\frac{n}{k})^{k}\leq\binom{n}{k}<(\frac{en}{k})^{k}$ and $k!\geq(k/e)^{k}$.
where $P=\mathrm{Bern}(p,p,\rho)$ and $Q=\mathrm{Bern}(p,p,0)$. Pick $\mathcal{Q}$ to be an auxiliary null model under which $G_{1}$ and $G_{2}$ are independent with the same marginal as $\mathcal{P}$. Then, $\mathcal{Q}_{G_{1},G_{2}}$ can be factorized into
for any $\pi:S\mapsto T$ with $|S|=|T|=m$. By Lemma 8 and (19), we obtain
$I(\pi^{*};G_{1},G_{2})\leq\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}})\leq\binom{m}{2}D(P\|Q)\leq 25\binom{m}{2}p^{2}\phi(\gamma).$ (52)
By Lemma 8 and (19), we obtain
$I(\pi^{*};G_{1},G_{2})\leq\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}})\leq\binom{m}{2}D(P\|Q)\leq\frac{1}{2}\binom{m}{2}\log\left(\frac{1}{1-\rho^{2}}\right).$ (53)
Appendix D Auxiliary results
@@
where the last inequality is due to the fact that $(1-x)\log(1-x)+x\geq 0$ for any $0<x<1$ and $0<\eta\leq\frac{\gamma}{4(1+\gamma)}<\frac{1}{4}$. Since $\eta\leq\frac{\gamma}{4(1+\gamma)}$ and $(1+\gamma)\log(1+\gamma)-\gamma\geq\frac{\gamma^{2}}{2(1+\gamma)}$, we obtain $\eta\gamma\leq\frac{\gamma^{2}}{4(1+\gamma)}\leq\frac{1}{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]$. Therefore,
@@
where $\mathrm{(a)}$ is because h(x)=h(1x)h(x)=h(1-x) and $h(x)\geq-x\log x$; $\mathrm{(b)}$ is because $\log\left(\frac{m}{k}\right)\geq\log 2$ when $k\leq\frac{m}{2}$; $\mathrm{(c)}$ is because $k\log\left(\frac{m}{k}\right)\geq\log m$ for $1\leq k\leq 2\log m$ when $m\geq 10$; $\mathrm{(d)}$ is because $2\cdot 2^{-2\log m}\leq\frac{2}{m}$. ∎
@@
$\overset{\mathrm{(b)}}{=}$ $~p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right],$
where $\mathrm{(a)}$ uses $\log(1+x)\leq x$ for any x>1x>-1; $\mathrm{(b)}$ follows from $\gamma=\frac{\rho(1-p)}{p}$. Since $\log(1+x)\geq\frac{x}{x+1}+\frac{x^{2}}{2(x+1)^{2}}$ for any $x\geq 0$, we obtain that $p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]\geq\frac{p^{2}\gamma^{2}}{2(\gamma+1)}$. When $\gamma<3$, since $\frac{p^{2}\gamma^{2}}{2(\gamma+1)}\geq\frac{p^{2}\gamma^{2}}{8}=\frac{\rho^{2}\cdot 3(1-p)^{2}}{24}\geq\frac{\rho^{2}\left[2p(1-p)+p^{2}\right]}{24}$ for $0<p\leq\frac{1}{2}$, we obtain that